OpenStudy (anonymous):

the jobless rate in a small town in nevada increased from 8.7% in 2007 to 12.3% in 2009 predict the jobless rate in 2010

OpenStudy (amistre64):

do we assume exponential or linear?

OpenStudy (a_clan):

cannot be determined

OpenStudy (amistre64):

linear prediction can be :)

OpenStudy (anonymous):

i don't understand this one can you explain

OpenStudy (amistre64):

2009,12.3 - 2007, 8.7 -------------- 2, 3.7 ; slope = 3.7/2

OpenStudy (amistre64):

y = (3.7/2)x -(3.7(2009)/2) + 12.3 y = 1.85x -3704.35 maybe

OpenStudy (amistre64):

my 3.7 should be 3.6 :)

OpenStudy (amistre64):

14.1 should be the answer eventually

OpenStudy (amistre64):

y = 1.8x - 3603.9 is the linear approximation